In order for you to be indifferent between the after tax returns on a corporate bond paying 8.0% and a tax-exempt municipal bond paying 6.0%, what would your tax bracket need to be?

A. 25%

B. 75%

C. 28%

D. 15%

E. Cannot tell from the information given
Answer #1

Return on corporate bond *(1-tax rate)=municipal bond yield

8*(1-tax rate)=6

1-tax rate=(6/8)

tax rate=1-(6/8)

which is equal to

=25%